What is the angle of deviation when the angle of incidence is 10, 20, 30, 40, 50, 60, 70
1
Expert's answer
2020-08-14T16:11:30-0400

Angle of deviation is depending on the reflection and refraction.

If reflection is getting happen, then angle of incidence = angle of reflection,

Hence deviation "=180-2i"

Here i is the angle of incidence.

If angle of incidence are 10, 20, 30, 40, 50, 60 and 70 then angle of deviation will be 160, 140, 120, 100, 80, 60 and 40.


For the refraction, the value of refractive index must be given in the question, but it is not mentioned in the question, so it can not be determined.

"\Rightarrow \frac{\mu_2}{\mu_1} = \frac{\sin i}{\sin r}"

Here, "\mu_2" and "\mu_1" are representing the refractive index of the medium, form the this we can calculate the angle of refraction then we can conclude the deviation.